# Optimizing compilation of error correction codes for $2\times N$ quantum dot arrays and its NP-hardness

### v1

@@ IV.1 Special Case of Qubit Re-indexing
As we were doing in the naive syndrome-extraction circuit case, we again only re-index ancillary qubits, however, whereas before we were doing so merely as a simplification done in order to make progress on the problem despite it not always finding the most optimal indexing of qubits, in the case of Shor-syndrome re-indexing, we conjecture that only re-indexing the ancillary qubits will always be sufficient for finding an optimal solution. We now note how the structure of Shor-syndrome circuits causes our problem definition to change. Because all primitive sets become size 1 (i.e. an ancillary qubit interacts with exactly one data qubit, ensuring fault-tolerance), instead of finding an assignment from integers to sets, we now only need to find an assignment between integers and integers. More formally, making use of all sets being singletons, we can simplify the problem formulation from the naive syndrome case to:
@@
Just as in the naive syndrome case, the minimization of $|\mathcal{G}|_{0}$ corresponds to the minimization of the number of required number of shuttles, and the $\pi$ we find in this process describes how we should reorder our ancillary qubits. Just like before, doing so also increases efficiency and decreases the amount of time required per syndrome extracting cycle. Refer to Fig. 6 for an example.
Figure 6: Feasible solution to re-indexing of Shor-style syndrome extraction for the Steane code. On the right is our set $\mathcal{S}$, derived from the Shor-style syndrome extraction circuit for the Steane code. On the left we have a non-optimal ordering. Here we are placing the numbers from the circle (each number corresponding to a gate in the syndrome circuit) onto this “ladder”, with the goal being to minimize the number of different output values obtaining by the addition of the placed value with the index it was placed in. Each output value corresponds to the $\delta$ of the placed gate after re-indexing. The ordering in this figure results in 7 different output (or $\delta$) values, although an optimal solution of 6 values exists, namely Fig. 8.
Figure 7: Circuit before and after Specialized Shor Syndrome Compilation. Here we have the syndrome extracting two qubit gates for measuring syndromes of the Steane code with Shor-style syndrome extraction before and after our compilation algorithm. Consecutive gates that are in a color-coded group together can run concurrently on the $2\times N$ architecture (repeated colors have no relation to each other). Furthermore, each grouping corresponds to a particular value of $\delta$. a.) Here is the uncompiled (yet gate shuffled) circuit. There are 14 values for delta in this circuit and therefore 14 shuttles are require to execute this. b.) Here is the circuit after our Specialized Shor Syndrome Compilation algorithm. Notice that now there are now only 6 values of $\delta$, and therefore this circuit will only require 6 shuttles to execute.
This new formulation allows us to make a few claims about the hypothetical optimal solution structure, which we can leverage in the design of our algorithm. First we notice that if a value $v$ occurs $f$ times in $\mathcal{S}$, then $\min_{\pi}|\mathcal{G}|_{0}\geq f$. This follows simply from us considering a sub-multiset $U:=\{v,v,...v\}\subseteq\mathcal{S}$, and $U$ contains $f$ elements. Now, if we add to each element of $U$ a value from $\{1,2,...,s\}$ without repetition, and call the resulting multiset $U^{\prime}$, then $U^{\prime}$ will contain exactly $f$ unique elements, regardless of how we choose elements from $\{1,...,s\}$.
@@
[…]
@@
[…]
@@
Depending on implementation, one may have some chains with leading or trailing $-1$’s. If desired, remove those.
@@
Figure 8: Optimal solution to Shor re-indexing on Steane code. Using Algorithms 1 and 2, we can find an optimal reordering of qubits for the Steane code’s Shor-style syndrome extraction circuit. On the right, we group the elements into six chains, the maximum column weight of the code. On the left, we show how we can pack these without breaking these chains up, giving us an optimal result of six shuttles. Each output value from mapping values to the ladder corresponds to a particular configuration of the $2\times n$ architecture.
@@
A sorted list of numbers whose values are the primitive sets, and whose ordering is the order that when $[1,...s]$ is added, results in an array with as ideally as few distinct values as possible.
@@
V Simulation Results
Through simulation, we show that our methods of circuit compilation can be used to reduce the threshold of various error correcting codes. We model Pauli noise with highly efficient Pauli-Frame simulation techniques [11, 12] implemented in Julia [13]. We simulate at least 20,000 samples for each set of noise parameter values (each data point in our plots) in order to calculate the logical error rate. The simulated circuits contain the following steps:
  • Logical operator measurement In order to have something to compare our after-decoding results to, we measure the logical X or Z operators at the end each syndrome circuit, as if it were an additional parity check. More specifically, we cannot measure both operators in the same circuit, so we need to run two simulations in order to obtain both the logical X and Z error rates.
In our plots, we vary $p_{\text{mem}}$ along the horizontal axis and use constant values for other types of error. Specifically, we use a value of $p_{\text{shuttle}}=0.0001$ from [4], a potentially near-term achievable value of $p_{\text{gate}}=0.9995$, and $p_{\text{wait}}=1-e^{\frac{-(2\mu\text{s}+12.5\mu\text{s})}{28\text{ms}}}$. The shuttle time was said to be $2\mu$s ramp and the fastest time between shuttles was stated to be $12.5\mu$s in [4]. $28$ms was the T2 time reported in [16] (recently great effort has gone into improving to these times [3]). These near-term parameters are represented by the triangles in Fig. 9. Notice that these curves have error floors due to constant error terms that do not depend on $p_{\text{mem}}$. However, the curves plotted by circles correspond to scaling $p_{\text{wait}}$, $p_{\text{shuttle}}$, and $p_{\text{gate}}$ by $p_{\text{mem}}$, while keeping ratios constant, creating curves with pseudo-thresholds, as one may be used to.
A table of shuttle reductions for various codes and levels of compilation is shown in Table 1 for naive syndrome circuits and in Table 2 for Shor-style syndrome extracting circuits. Table 3 displays the number of required shuttles when considering the $X$ and $Z$ checks separately for Shor-style syndrome extraction. For plots of physical against logical error rates for different levels of compilation for codes other than the $3\times 3$ toric code, refer to the Appendix.
Code Uncompiled Gate Shuffled Ancillary Re-indexing Ancillary + Data Re-indexing
Steane7 24 11 9 7
Shor9 22 10 6 6
Perfect5: 16 6 6 6
Toric 3x3: 64 12 21 16
Toric 6x6: 280 12 45 43
Toric 100x100: 79,992 12 797 607
$[[360,60,d]]$ qLDPC code A: 2034 511 363 361
$[[360,60,d]]$ qLDPC code B: 2034 515 370 365
$[[144,12,12]]$ BB code: 864 24 110 122
$[[360,12,24]]$ BB code: 2160 24 241 287
Table 1: Different compilations of the naive syndrome circuit for extracting all stabilizers.. After generating the circuit for naively measuring the syndromes of all stabilizers for the given code, we performed four levels of circuit optimization: Uncompiled: the gate order was decided in the order that stabilizers are arbitrarily written, Gate Shuffled: the gate order was sorted by each gate’s $\delta$ as described in Section II.1, Ancillary Re-indexing (AHR): the ancillary qubits were re-indexed by the heuristics outlined in Section II.2, followed then by gate shuffling, and Ancillary + Data Re-indexing: after performing ancillary qubit re-indexing (AHR), we then fix the ancillary qubits instead and perform the same AHR heuristics on the data qubits. The two qLDPC codes (A and B) are small pedagogical instances of expander graph codes [17]. BB codes are bi-variate bicycle codes as used in IBM’s recent paper [18]
Code Uncompiled Gate Shuffled AHR SSSC
Steane7 14 14 7 6
Shor9 6 6 6 5
Perfect5: 6 6 4 4
Toric 3x3: 53 42 9 8
Toric 6x6: 221 173 9 8
Toric 100x100: 60,193 40,262 9 9
$[[360,60,d]]$ qLDPC code A: 1536 1069 272 122
$[[360,60,d]]$ qLDPC code B: 1608 1120 260 119
$[[144,12,12]]$ BB code: 744 446 6 6
$[[360,12,24]]$ BB code: 1860 1098 6 6
Table 2: Different compilations of the fault tolerant Shor-style syndrome circuit for extracting all stabilizers.. As in Table 1, we performed multiple levels of circuit optimization, however the provided circuit to be optimized is the fault tolerant Shor-style syndrome extracting circuit, as described in Section IV. We conjecture that ancillary + data qubit re-indexing doesn’t offer any advantage to Shor syndrome circuits so that column is not present, however in its stead we have the algorithm specialized for these types of circuits, also from Section IV, Shor Syndrome Specialized Compilation (SSSC).
Code Uncompiled Gate Shuffled AHR SSSC
Steane7 7 (7) 7 (7) 3 (3) 3 (3)
Shor9 2 (4) 2 (4) 2 (4) 2 (4)
Toric 3x3: 27 (26) 21 (21) 4 (5) 4 (4)
Toric 6x6: 111 (110) 87 (86) 4 (5) 4 (5)
Toric 100x100: 30,097 (30,096) 20,131 (20,131) 4 (5) 4 (5)
$[[360,60,d]]$ qLDPC code A: 756 (780) 547 (522) 195 (106) 58 (54)
$[[360,60,d]]$ qLDPC code B: 828 (780) 598 (522) 196 (106) 64 (54)
$[[144,12,12]]$ BB code: 372 (372) 229 (229) 3 (3) 3 (3)
$[[360,12,24]]$ BB code: 930 (930) 549 (549) 3 (3) 3 (3)
Table 3: Different compilations of the fault tolerant Shor-style syndrome circuit for extracting all X (Z) stabilizers on CSS codes. Similar to Table 2, except here we compile a circuit that extracts only X (Z) stabilizers. The reason for doing this would be to completely ensure commutativity of gate operations on any CSS code.

### v2

@@ IV.1 Special Case of Qubit Re-indexing
As we were doing in the naive syndrome-extraction circuit case, we again only re-index ancillary qubits, so that the data qubits retain the same labeling between the $X$ and $Z$ circuits. However unlike before, we conjecture that in the case of Shor syndrome circuits, re-indexing the data qubits will never lead to an advantage in shuttle minimization.
We now note how the structure of Shor-syndrome circuits causes our problem definition to change. Because all primitive sets become size 1 (i.e. an ancillary qubit interacts with exactly one data qubit, ensuring fault-tolerance), instead of finding an assignment from integers to sets, we now only need to find an assignment between integers and integers. More formally, making use of all sets being singletons, we can simplify the problem formulation from the naive syndrome case to:
@@
Just as in the naive syndrome case, the minimization of $|\mathcal{G}|_{0}$ corresponds to the minimization of the number of required number of shuttles, and the $\pi$ we find in this process describes how we should reorder our ancillary qubits. As before, doing so also increases efficiency and decreases the amount of time required per syndrome extracting cycle. Refer to Fig. 6 for an example.
Figure 6: Feasible solution to re-indexing of Shor-style syndrome extraction for the Steane code. On the right is our set $\mathcal{S}$, derived from the Shor-style syndrome extraction circuit for the Steane code (X and Z checks are compiled together to create a more pedgocical example). On the left we have a non-optimal ordering. Here we are placing the numbers from the circle (each number corresponding to a gate in the syndrome circuit) onto this “ladder”, with the goal being to minimize the number of different output values obtaining by the addition of the placed value with the index it was placed in. Each output value corresponds to the $\delta$ of the placed gate after re-indexing. The ordering in this figure results in 7 different output (or $\delta$) values, although an optimal solution of 6 values exists, namely Fig. 7.
Figure 7: Optimal solution to Shor re-indexing on Steane code. Using Algorithms 1 and 2, we can find an optimal reordering of qubits for the Steane code’s Shor-style syndrome extraction circuit. On the right, we group the elements into six chains, the maximum column weight of the code. On the left, we show how we can pack these without breaking these chains up, giving us an optimal result of six shuttles. Each output value from mapping values to the ladder corresponds to a particular configuration of the $2\times n$ architecture. Note: as with Fig. 6, the $X$ and $Z$ checks are presented together for pedagogical reasons.
This new formulation allows us to make a few claims about the hypothetical optimal solution structure, which we can leverage in the design of our algorithm. First, we notice that if a value $v$ occurs $f$ times in $\mathcal{S}$, then $\min_{\pi}|\mathcal{G}|_{0}\geq f$. This follows simply from us considering a sub-multiset $U:=\{v,v,...v\}\subseteq\mathcal{S}$, and $U$ contains $f$ elements. Now, if we add to each element of $U$ a value from $\{1,2,...,s\}$ without repetition, and call the resulting multiset $U^{\prime}$, then $U^{\prime}$ will contain exactly $f$ unique elements, regardless of how we choose elements from $\{1,...,s\}$.
@@
Initialize $Y$ to a 2D array of size $\text{num\_chains}\times\text{max\_value}$, each element is equal to 1-1.
@@
  $\text{index}\leftarrow$ index($x$) $\triangleright$ get index in $Y$ that corresponds to the value $x$
  $\text{placed}\leftarrow$ False
  for i in $1:\text{num\_chains}$ do
@@
Depending on implementation, one may have some chains with leading or trailing 1-1’s. If desired, remove those.
@@
A sorted list of numbers whose values are the primitive sets, and whose ordering is the order that when [1,s][1,...s] is added, results in an array with as ideally as few distinct values as possible.
@@
Figure 8: Steane code’s shor-style syndrome extraction circuit’s two qubit gates, both before and after Specialized Shor Syndrome Compilation. Consecutive gates that are in a color-coded group together can run concurrently on the $2\times n$ architecture (repeated colors have no relation to each other). Furthermore, each grouping corresponds to a particular value of $\delta$. a.) The uncompiled (yet gate shuffled) circuit. There are 14 values for delta in this circuit and therefore 14 shuttles are require to execute this. b.) The circuit after our Specialized Shor Syndrome Compilation algorithm. Notice that now there are now only 6 values of $\delta$, and therefore this circuit will only require 6 shuttles to execute.
V Compilation Results
Tables 1 and  2 host the number of shuttles required for extracting both the X and Z syndromes of various CSS codes on Shor and naive syndrome circuits respectively. Furthermore, in Appendix. .4, we show through gate-by-gate simulation how the logical error rate decreases with compilation on the toric code.
V.1 Codes Tested
All codes tested were generated with QuantumClifford.jl [14]. Specifically, following the QuantumClifford’s ECC API, stablizer tableau were obtained, and then the appropriate methods were called to isolate the X and Z checks, and then to generate the syndrome extracting circuit to be compiled.
  • 2-dimensional Toric and Surface codes: $[[18,2,3]]$, $[[18,2,3]]$, $[[50,2,4]]$ 2D toric codes [15] and $[[13,1,3]]$, $[[25,1,4]]$, $[[41,1,5]]$ 2D surface codes [16] with lattice sizes ranging from $3\times 3$ to $5\times 5$ were tested.
  • Shor’s [[9, 1, 3]] code [17]
  • Steane [[7, 1, 3]] code [18]
  • 2D Color codes: The triangular planar codes with 4.8.8 tiling, given by the instances $[[17,1,5]]$, $[[31,1,7]]$, and $[[71,1,11]]$, and those with 6.6.6 tiling, given by $[[19,1,5]]$, $[[37,1,7]]$, and $[[91,1,11]]$, are considered [19].
  • Lifted Product (LP) codes: A $882,24,d\leq 24]]$ code from Appendix B of [20], and a $175,19,\leq 10]]$ code Appendix A of [21].
  • Generalized Bicycle codes: The $[[48,6,8]]$ and $254,28,14\leq d\leq 20]]$ code instances from Appendix B of [22].
  • Two Block Group Alegbra (2BGA) codes: The $[[40,8,5]]$, $[[56,28,2]]$, and $[[70,8,10]]$ abelian 2BGA codes from Table I of [23], as well as $[[64,8,8]]$ non-abelian 2BGA code from Table II of the same work [23].
  • Multivariate Bicycle code: A weight-$6$ $[[48,4,6]]$ code from [24].
  • Bivariate Bicycle codes: The $[[144,12,12]]$, $360,12,\leq 24]]$, and $[[756,16,34]]$, as reported in Table 3 of [25].
  • Trivariate Tricycle codes: The $[[72,6,6]]$, $[[432,12,12]]$, $[[840,9,16]]$ and $[[1029,18,16]]$, as reported in Table V of [26].
  • La-Cross codes: The instances $[[100,4,5]]$, $[[117,9,4]]$, $[[296,16,7]]$, and $[[400,16,8]]$ from the $(n-k)^{2}+n^{2},k^{2},d]]$ La–Cross code family with open boundary conditions, together with the instance $[[98,18,4]]$ from the $2n^{2},2k^{2},d]]$ La–Cross code family with periodic boundary conditions, are reported in Fig. 3 of [27].
  • Quasi-cyclic Generalized Hypergraph Product code: We consider the code instance $[[882,48,16]]$, in which half of the columns have weight 3 and the remaining half have weight 5, as reported in [22]. Building on this construction, we generate a new instance with parameters $[[882,12,d]]$, where half of the columns have weight 3 and the other half have weight 7. This construction employs a sparse quasi–cyclic polynomial matrix $A$ over the quotient ring $R=\mathbb{F}_{2}[x]/(x^{\ell}-1)$, with each entry being either 0 or a monomial $x^{i}$ for $0\leq i<\ell$, where $\ell$ denotes the circulant size. Each polynomial entry corresponds to a binary $\ell\times\ell$ circulant block. To ensure sparsity, each row and column of $A$ contains at most $w$ nonzero entries, so that the resulting parity–check matrices are $w+\deg b(x)$–limited. A fixed polynomial $b(x)\in R$, referred to as the circulant polynomial, is used to construct the off-diagonal blocks. For the present code instance, we take $\ell=63$ and construct a matrix $A\in M_{n\times n}(R)$ with $n=7$, i.e., $A$ is a $7\times 7$ matrix over $R$:
    $\small A=\begin{pmatrix}x^{56}&x^{12}&x^{52}&x^{13}&x^{39}&x^{11}&x^{61}\\ x^{61}&x^{56}&x^{12}&x^{52}&x^{13}&x^{39}&x^{11}\\ x^{11}&x^{61}&x^{56}&x^{12}&x^{52}&x^{13}&x^{39}\\ x^{39}&x^{11}&x^{61}&x^{56}&x^{12}&x^{52}&x^{13}\\ x^{13}&x^{39}&x^{11}&x^{61}&x^{56}&x^{12}&x^{52}\\ x^{52}&x^{13}&x^{39}&x^{11}&x^{61}&x^{56}&x^{12}\\ x^{12}&x^{52}&x^{13}&x^{39}&x^{11}&x^{61}&x^{56}\end{pmatrix}b(x)=1+x+x^{6}\in R$
    The code is then defined via the block parity–check matrices
    $H_{X}=\begin{bmatrix}A,&b(x)I_{m}\end{bmatrix},\qquad H_{Z}=\begin{bmatrix}b(x)^{T}I_{n},&A^{*}\end{bmatrix}$
    where $A^{*}$ denotes the transpose of $A$ with each entry polynomial-reversed modulo $x^{\ell}-1$, and $I_{m},I_{n}$ are identity matrices over $R$ [22]
  • 3D and 4D Surface codes: The instances of the 3D surface code are $[[12,1,2]]$ for $L=2$ and $[[51,1,3]]$ for $L=3$, while the 4D surface code instances are $[[33,1,4]]$ for $L=2$ and $[[241,1,9]]$ for $L=3$. These codes are constructed via the hypergraph product of $(L-1)\times L$ repetition code chain complexes. The $[[33,1,4]]$ 4D surface code and the construction method for these codes are described in [28].
  • Generalized Toric codes on twisted tori: We consider the code instances $[[292,18,8]]$, $392,8,\leq 28]]$, and $396,8,\leq 36]]$, as summarized in Tables III and IV of [29].
  • Homological Product codes: We consider the code instances $[[117,9,4]]$ and $[[225,9,6]]$, as reported in Table III of [30].
  • Double Homological Product codes: We consider the code instances $[[241,1,9]]$ and $[[486,6,9]]$, as reported in Table I of [31].
V.2 Shor Compilation
The number of shuttling operations to carry out the fault-tolerant Shor syndrome extracting X and Z circuits for these codes under various levels of compilation is displayed in Table 1. The first number in the parenthesis corresponds to the value for the X check syndrome circuit, and the second value corresponds to the Z circuit. Levels of compilation are defined as follows:
  • Uncompiled: The number of shuttles to run the original circuit provided by QuantumClifford.jl, reading the circuit from left to right, shuttling whenever is needed to run the next gate.
  • GateShuffled: The number of shuttles required after $\delta$ sorting the circuit as discussed in Section II.1.
  • AHR: The number of shuttles after reindexing the ancilla by the simple sorting heuristics discussed in Section III.
  • SSSC: The number of shuttles after reindexing the ancilla by our specialized Shor syndrome circuit algorithm, discussed in Section IV.
  • NumChains & Blanks: The “NumChains” value corresponds to the minimum number of shuttles possible, achievable by inserting “blanks” into the ancilla array. The “Blanks” value is then a naive upper bound on the number of blanks required to achieve this. It was calculated by counting the number of gaps in the chains. Chain were introduced in Section IV, and the idea of adding blanks to achieve the minimum number of possible shuttles is discussed further in Section VI
In Table 1, we can see that our techniques find quite good shuttling schedules, trivially so on codes with regular column weight (and so SSSC does not add anything). For codes with non-regular column weight, with only some exception, SSSC does further reduce the number of shuttles compared to our heuristic sorting methods. Furthermore notice that we need not use the same reindexing technique for both the $X$ and $Z$ circuits. For example on the $[[117,9,4]]$ homological product code tested, we could use AHR to compile the X circuit to run in 12 shuttles, while using SSSC to compile the Z circuit to run in 7 shuttles. This is also exemplified in the $[[117,9,4]]$ La-Cross code, which can run with 5 shuttles on the X circuit after compiling via AHR, and then 5 shuttles on the Z circuit after compiling it with SSSC.
Another thing to notice is that since these methods are taking as input a circuit algorithmically generated via its stabilizer tableau, codes like the 2D toric code compile to different shuttle numbers for its X and Z circuits. This is simply due to the stabilizer tableaux for its X and Z checks being literally different due to how they are generated in QuantumClifford.jl [14]. The manifestation of the 2D toric code’s stabilizer generators form a column-irregular matrix, at least as defined in QuantumClifford.jl, and so optimal compilation is not promised. An interesting avenue for future work might be on techniques that use Gaussian elimination on the stabilizer tableau to work in tandem with our methods which merely permute it.
V.3 Naive Compilation
The number of shuttling operations to carry out the naive syndrome extracting circuits for these codes are displayed in Table 2. The definitions of the types of compilation are the same as for Shor syndrome extraction. Naturally, there are fewer, as we developed some specialized methods for Shor syndrome extraction. For most of the codes tested, AHR does not reduce the number of shuttles for their naive syndrome circuit. This is due to naive compilation being a much harder problem than Shor syndrome compilation, which we show is NP-hard in Appendix .2. The AHR methods are after all heuristics and need not be used if the original ordering of ancilla is better. Development of algorithms to improve performance of more general or other special classes of syndrome extraction circuits (flag qubit circuits are perhaps more interesting than these naive circuits), we leave to future research.
Code N K D Uncompiled GateShuffled AHR SSSC NumChains Blanks
2DTORIC 18 2 3 (27, 26) (21, 21) (4, 5) (4, 4) (2, 2) (2, 3)
2DTORIC 32 2 4 (49, 48) (35, 35) (4, 5) (4, 5) (2, 2) (2, 3)
2DTORIC 50 2 5 (77, 76) (62, 61) (4, 5) (4, 5) (2, 2) (2, 3)
2DSURFACE 13 1 3 (18, 14) (14, 13) (3, 5) (3, 4) (2, 2) (3, 5)
2DSURFACE 25 1 4 (36, 30) (26, 25) (3, 6) (3, 4) (2, 2) (4, 7)
2DSURFACE 41 1 5 (60, 52) (48, 34) (3, 7) (3, 5) (2, 2) (5, 9)
SHOR 9 1 3 (2, 4) (2, 4) (2, 4) (2, 4) (2, 2) (0, 4)
STEANE 7 1 3 (7, 7) (7, 7) (3, 3) (3, 3) (3, 3) (1, 1)
2DCOLOR4.8.8 17 1 5 (19, 19) (15, 15) (7, 7) (4, 4) (3, 3) (7, 7)
2DCOLOR4.8.8 31 1 7 (38, 38) (27, 27) (9, 9) (7, 7) (3, 3) (10, 10)
2DCOLOR4.8.8 71 1 11 (94, 94) (67, 67) (13, 13) (9, 9) (3, 3) (18, 18)
2DCOLOR6.6.6 19 1 5 (22, 22) (17, 17) (7, 7) (6, 6) (3, 3) (5, 5)
2DCOLOR6.6.6 37 1 7 (48, 48) (32, 32) (10, 10) (7, 7) (3, 3) (9, 9)
2DCOLOR6.6.6 91 1 11 (127, 127) (82, 82) (16, 16) (12, 12) (3, 3) (17, 17)
LIFTEDPRODUCT 882 24 $d\leq$ 24 (2212, 2212) (1820, 1820) (3, 3) (3, 3) (3, 3) (0, 0)
LIFTEDPRODUCT 175 19 $d\leq$ 10 (582, 584) (395, 377) (4, 4) (4, 4) (4, 4) (0, 0)
GENERALIZEDBICYCLE 48 6 8 (192, 192) (139, 139) (4, 4) (4, 4) (4, 4) (0, 0)
GENERALIZEDBICYCLE 254 28 $14\leq d\leq 20$ (1144, 1144) (827, 827) (5, 5) (5, 5) (5, 5) (0, 0)
TWOBLOCKGROUPALGEBRA 70 8 10 (212, 212) (145, 145) (4, 4) (4, 4) (4, 4) (0, 0)
TWOBLOCKGROUPALGEBRA 40 8 5 (121, 121) (74, 74) (6, 6) (6, 6) (6, 6) (0, 0)
TWOBLOCKGROUPALGEBRA 56 28 2 (142, 142) (136, 136) (6, 6) (6, 6) (6, 6) (0, 0)
TWOBLOCKGROUPALGEBRA 64 8 8 (224, 224) (159, 168) (6, 6) (6, 6) (6, 6) (0, 0)
MULTIVARIATEBICYCLE 48 4 6 (144, 144) (93, 93) (4, 4) (4, 4) (4, 4) (0, 0)
LA-CROSS 100 4 5 (192, 156) (135, 99) (5, 19) (12, 15) (3, 3) (24, 45)
LA-CROSS 98 18 4 (252, 252) (201, 201) (3, 3) (3, 3) (3, 3) (0, 0)
LA-CROSS 117 9 4 (240, 216) (180, 160) (5, 21) (12, 5) (3, 3) (45, 77)
LA-CROSS 296 16 7 (630, 580) (461, 458) (5, 31) (46, 27) (3, 3) (98, 163)
LA-CROSS 400 16 8 (876, 816) (618, 605) (5, 35) (52, 39) (3, 3) (112, 187)
BIVARIATEBICYCLE 144 12 12 (372, 372) (229, 229) (3, 3) (3, 3) (3, 3) (0, 0)
BIVARIATEBICYCLE 360 12 $d\leq$24 (930, 930) (549, 549) (3, 3) (3, 3) (3, 3) (0, 0)
BIVARIATEBICYCLE 756 16 $d\leq$34 (2268, 2268) (1442, 1442) (3, 3) (3, 3) (3, 3) (0, 0)
TRIVARIATETRICYCLE 72 6 6 (207, 415) (148, 315) (3, 6) (3, 6) (3, 6) (0, 0)
TRIVARIATETRICYCLE 432 12 12 (1266, 2532) (838, 1722) (3, 6) (3, 6) (3, 6) (0, 0)
TRIVARIATETRICYCLE 840 9 16 (2520, 5040) (1642, 3214) (3, 6) (3, 6) (3, 6) (0, 0)
TRIVARIATETRICYCLE 1029 18 16 (3087, 6174) (1890, 4429) (3, 6) (3, 6) (3, 6) (0, 0)
3DSURFACE 12 1 2 (9, 23) (9, 16) (3, 4) (3, 4) (2, 3) (4, 4)
3DSURFACE 51 1 3 (66, 138) (46, 90) (13, 10) (10, 8) (2, 4) (17, 31)
4DSURFACE 33 1 4 (64, 75) (49, 54) (8, 8) (7, 7) (4, 4) (16, 16)
4DSURFACE 241 1 9 (656, 715) (419, 461) (67, 37) (44, 27) (4, 4) (152, 196)
GENERALIZEDTORIC 292 18 8 (730, 730) (504, 504) (3, 3) (3, 3) (3, 3) (0, 0)
GENERALIZEDTORIC 392 6 $d\leq$28 (280, 280) (196, 196) (3, 3) (3, 3) (3, 3) (0, 0)
GENERALIZEDTORIC 396 8 $d\leq$36 (991, 991) (583, 583) (3, 3) (3, 3) (3, 3) (0, 0)
HOMOLOGICALPRODUCT 117 9 4 (225, 225) (155, 162) (12, 8) (22,7) (3, 3) (27, 42)
HOMOLOGICALPRODUCT 225 9 6 (618, 630) (432, 447) (27, 14) (28, 11) (4, 4) (72, 120)
DBLEHOMOLOGICALPRODUCT 486 6 9 (1782, 1782) (1204, 1204) (4, 4) (4, 4) (4, 4) (0, 0)
QUASICYCLIC_GHP 882 48 16 (3093, 3093) (2151, 2151) (5, 5) (5, 5) (5, 5) (0, 0)
QUASICYCLIC_GHP 882 12 (3976, 3976) (2829, 2829) (7, 7) (7, 7) (7, 7) (0, 0)
Table 1: Compilation numbers for Shor-style syndrome circuits for extracting X, Z stabilizers separately on CSS codes. As described in Section. V.2, the different columns correspond to the required numbers of shuttles corresponding to different levels of compilation (the number of shuttles is listed for $X$ circuits first and then $Z$). However “NumChains” corresponds to the smallest number of shuttles possible, obtainable by introducing “blanks” into the ancilla array. A very naive upper bound on the number of blanks required for perfect compilation is presented in the “Blanks” column, rather than a number of shuttles. Refer to Section VI for more on blanks and perfect compilation. Quantum error correcting code parameters: N is the number of physical qubits, K the number of logical qubits, and true code distance $\textbf{D}=\min(d_{X},d_{Z})$ is defined as the minimum weight of a non-trivial logical Pauli operator. Distances were calculated using the mixed-integer programming method of [19].
Code N K D Uncompiled GateShuffled AHR
2DTORIC 18 2 3 (32, 32) (6, 6) (11, 10)
2DTORIC 32 2 4 (60, 60) (6, 6) (12, 11)
2DTORIC 50 2 5 (96, 96) (6, 6) (12, 11)
2DSURFACE 13 1 3 (20, 20) (5, 6) (5, 6)
2DSURFACE 25 1 4 (42, 42) (6, 7) (6, 7)
2DSURFACE 41 1 5 (72, 72) (7, 8) (7, 8)
SHOR 9 1 3 (12, 10) (8, 4) (8, 4)
STEANE 7 1 3 (12, 12) (8, 8) (7, 7)
2DCOLOR4.8.8 17 1 5 (36, 36) (13, 13) (13, 13)
2DCOLOR4.8.8 31 1 7 (72, 72) (21, 21) (20, 20)
2DCOLOR4.8.8 71 1 11 (180, 180) (43, 43) (38, 38)
2DCOLOR6.6.6 19 1 5 (42, 42) (13, 13) (11, 11)
2DCOLOR6.6.6 37 1 7 (90, 90) (23, 23) (20, 20)
2DCOLOR6.6.6 91 1 11 (240, 240) (52, 52) (47, 47)
LIFTEDPRODUCT 882 24 d $\leq$ 24 (2646, 2646) (13, 13) (81, 32)
LIFTEDPRODUCT 175 19 d $\leq$ 10 (588, 588) (46, 53) (103, 55)
GENERALIZEDBICYCLE 48 6 8 (192, 192) (13, 13) (38, 36)
GENERALIZEDBICYCLE 254 28 14 $\leq$ d $\leq$ 20 (1270, 1270) (18, 18) (67, 67)
TWOBLOCKGROUPALGEBRA 70 8 10 (280, 280) (14, 14) (42, 40)
TWOBLOCKGROUPALGEBRA 40 8 5 (160, 160) (17, 17) (41, 31)
TWOBLOCKGROUPALGEBRA 56 28 2 (224, 224) (15, 15) (24, 15)
TWOBLOCKGROUPALGEBRA 64 8 8 (256, 256) (55, 55) (67, 66)
MULTIVARIATEBICYCLE 48 4 6 (144, 144) (16, 16) (35, 23)
LA-CROSS 100 4 5 (252, 252) (16, 20) (16, 20)
LA-CROSS 98 18 4 (294, 294) (10, 10) (37, 28)
LA-CROSS 117 9 4 (270, 270) (16, 22) (16, 22)
LA-CROSS 296 16 7 (720, 720) (24, 32) (24, 32)
LA-CROSS 400 16 8 (1008, 1008) (28, 36) (28, 36)
BIVARIATEBICYCLE 144 12 12 (432, 432) (12, 12) (21, 68)
BIVARIATEBICYCLE 360 12 d $\leq$ 24 (1080, 1080) (12, 12) (43, 186)
BIVARIATEBICYCLE 756 16 d $\leq$ 34 (2268, 2268) (19, 19) (39, 58)
TRIVARIATETRICYCLE 72 6 6 (216, 432) (28, 53) (55, 73)
TRIVARIATETRICYCLE 432 12 12 (1296, 2592) (46, 90) (86, 217)
TRIVARIATETRICYCLE 840 9 16 (2520, 5040) (43, 84) (98, 254)
TRIVARIATETRICYCLE 1029 18 16 (3087, 6174) (39,76) (107, 390)
3DSURFACE 12 1 2 (16, 27) (10, 9) (9, 9)
3DSURFACE 51 1 3 (84, 150) (16, 22) (16, 33)
4DSURFACE 33 1 4 (83, 84) (23, 18) (23, 18)
4DSURFACE 241 1 9 (758, 759) (72, 57) (79, 83)
GENERALIZEDTORIC 292 18 8 (876, 876) (230, 230) (189, 201)
GENERALIZEDTORIC 392 8 $d\leq$28 (336, 336) (95, 95) (95, 93)
GENERALIZEDTORIC 396 8 $d\leq$36 (1188, 1188) (312, 312) (251, 243)
HOMOLOGICALPRODUCT 117 9 4 (225, 225) (28, 22) (43, 52)
HOMOLOGICALPRODUCT 225 9 6 (630, 630) (38, 32) (49, 114)
DBLEHOMOLOGICALPRODUCT 486 6 9 (1944, 1944) (83, 83) (259, 91)
QUASICYCLIC_GHP 882 48 16 (3528, 3528) (19, 19) (102, 49)
QUASICYCLIC_GHP 882 12 (4410, 4410) (31, 31) (145, 126)
Table 2: Different compilations of the naive syndrome circuit for extracting all stabilizers. Same as Table 1, except for naive syndrome circuits. Naturally there are fewer columns, as we developed additional methods specifically for Shor-syndrome extraction.